A committee of 6 people is to be chosen from a group consisting of 7 men and 8 women. If the committee must consist of at least 3 women and at least 2 men, how many different committees are possible?
Answer by sudhanshu_kmr(1152)


Possible ways are :
( 3 women, 3 men) no. of ways = 8C3 * 7C3 = 56 * 35 = 1960
and
(4 women, 2 men) no. of ways = 8C4 * 7C2 = 70 * 21 = 1470

total no. of possible ways = 1960 + 1470 = 3430